OpenStudy (anonymous):

Write A Equation for the line m=5/6 and they y intercept is (0,12)

pooja195 (pooja195):

In slope intercept form?

pooja195 (pooja195):

Or point slope?

OpenStudy (anonymous):

Im Guessing Slop Intercept But They Didnt Say

pooja195 (pooja195):

I think so too... y=mx+b m=slope (x,y)

pooja195 (pooja195):

b= y intercept

OpenStudy (anonymous):

ok.

pooja195 (pooja195):

so what would the equation be?

OpenStudy (anonymous):

y=5/6x+0,2

pooja195 (pooja195):

You have it somewhat correct....(the slope is correct) y=5/6+____ 12 is the y intercept because (x,y) <---2nd coordinate is always y so ib b=12 what would the equation be? :)

pooja195 (pooja195):

Any ideas?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=5/6x+12

pooja195 (pooja195):

Yes! :D Well done!! Good job. :)

OpenStudy (anonymous):

Thanks

pooja195 (pooja195):

No probs :) let me know if you need more help.

OpenStudy (anonymous):

Mind Helping Me With This

pooja195 (pooja195):

I can try.

OpenStudy (anonymous):

m=0 and the y is 0,-2

pooja195 (pooja195):

same thing.... y=mx+b -2=b 0=slope

OpenStudy (anonymous):

Oh So Basically Just Plug in

pooja195 (pooja195):

Yep

OpenStudy (anonymous):

ok look what about this

OpenStudy (anonymous):

write each question in slope intercept form for y-intercept of each line 5x+y=4

pooja195 (pooja195):

subtract 5x from both sides so you have an equation in slope intercept form

OpenStudy (anonymous):

y=9